OpenStudy (anonymous):

the fourth and fifth terms of an arithmetic sequence are 13 and 15. find the 20th term

OpenStudy (anonymous):

let first term be a and common difference be d. a+3d=13 a+4d=15 subtracting second eqn from the first, we get d=2. substituting in 1st eqn, we get a+6=13 a=7. therefore 20th term is a+19d=7+38=45

OpenStudy (anonymous):

t4= a+3d=13 t5=a+ 4d=15 solving d=2 and a =7 t20 = a +19d=7+19*2=7+38=45